Good day,

When a solution of KOH is neutralized by hydrochloric acid, the solution gets hot.

Calculate the change in heat of the solution (in J).


Mass of combined solution = 42.0 g

Temperature change of the solution = 13.4 K

Specific heat capacity of the solution = 4.18 J/g·K


Please Give your answer to three significant figures.


Thanks


1
Expert's answer
2020-08-21T04:11:09-0400

Mass of solution "(m)= 42.0\ g"

Temperature change "(\triangle T)=13.4\ K"

Specific heat capacity"(C)=4.18J\/g.K"

Heat change will be "mC\triangle T=42.0\times 4.18\times 13.4=2352.504\ J"

"=2.35\times 10^3 J" (in 3 significant figures)